40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 17 22:51:50 2013, Koji, Update, Green Locking, ALS Y performance with the new whitening filter ALS_Y_130717.pdfALS_Y2_StripTool.png
    Reply  Thu Jul 18 01:10:00 2013, kiwamu, Update, Green Locking, ALS Y performance with the new whitening filter 
    Reply  Fri Jul 19 12:00:12 2013, manasa, Update, Green Locking, ALS Y performance with the new whitening filter 
Message ID: 8865     Entry time: Wed Jul 17 22:51:50 2013     Reply to this: 8866   8878
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: ALS Y performance with the new whitening filter 

[Manasa Koji]

Summary:
The new whitening filters improved the out-of-loop ALS stability of the Y arm down to 300Hz (20pm_rms in displacement).


- After modifying the whitening filters, the out-of-loop stability of the arms were tested with the IR PDH signals.

- The X arm showed non-stationarity and it made the ALS servo frequenctly fell out of lock.

- For now we decided to use the Y arm for the PRMI+one arm trial.

- The performance of the ALS was tested with several measurements. (attachment 1)

Cyan: Stability of the beatnote frequency with the MC and the arm freely running. The RMS of the day was ~6MHz.

Blue: Sensing limit of the beat box was tested by giving a signal from Marconi. The same amplitude as the X arm beat was given as the test signal.
This yielded the DC output of ~1200 counts.

Green: Out-of-loop estimation of the beatbox performance. This beat note stability was measured by controlling the arm with the IR PDH signal.
Assuming the PDH signal has better SNR than the beat signal, this gives us the out-of-loop estimation of the stability below 150Hz, which is the
unity gain frequency of the ALS loop.
Above 150Hz the loop does not force this noise to the suspension. Just the noise is injected via a residual control gain (<1).

Black: In-loop evaluation of the ALS loop. This becomes the left over noise for the true stability of the arm (for the IR beam).

Red: The arm was brought to the IR resonance using the ALS offset. The out-of-loop stability was evaluated by the IR PDH signal.
This indeed agreed with the evaluation with the other out-of-loop evaluation above (Green) below 150Hz.


Attachment 2 shows the time series data to show how the arm is brought to the resonance.
1 count of the offset corresponds to ~20kHz. So the arm started from 200kHz away from the resonance
and brought to the middle of the resonance.

(Manasa downloaded the 2k sampled data so that we can use this for presentations.)

Attachment 1: ALS_Y_130717.pdf  367 kB  | Hide | Hide all | Show all
ALS_Y_130717.pdf
Attachment 2: ALS_Y2_StripTool.png  17 kB  | Show | Hide all | Show all
ELOG V3.1.3-